OpenStudy (anonymous):

Which is the graph of y = 2.5x?

OpenStudy (campbell_st):

is the question 1. \[y = 2.5x\] or 2. \[y = 2.5^x\] either way use this site to graph the equation and compare it to your choices https://www.desmos.com/calculator

OpenStudy (anonymous):

Thank you do much! I'll fan and award you IF it's correct haha. @campbell_st

OpenStudy (anonymous):

Could you help me with another one?

OpenStudy (anonymous):

@campbell_st ?

OpenStudy (mathstudent55):

1. It's a straight line. That means the graph can be written in the form \(y = mx + b\) 2. Form the graph, you see it goes through point (0, 0), the origin, right?

OpenStudy (anonymous):

Um yeah.

OpenStudy (mathstudent55):

That means, the b in y = mx + b is zero. We have \(y = mx + 0\) or simply \(y = mx\) Now we need to find the slope, m.

OpenStudy (mathstudent55):

The line goes through (0, 0). Start at (0, 0), and going up to the right, can you find a point that is easy to read? Look for a point exactly on the intersection of rtwo grid lines.

OpenStudy (anonymous):

6,5?

OpenStudy (mathstudent55):

You're on the right track, but you mean (5, 6) Remember, you read x (left and right) first, then y (up and down).

OpenStudy (anonymous):

Sorry! Haha I suck so hard at math.

OpenStudy (anonymous):

I'm still a little confused @mathstudent55

OpenStudy (mathstudent55):

Ok, we have point (0, 0) and point (5, 6). The slope is a ratio. A ratio is usually expressed as a fraction. The slope is the ratio of change in y to the change in x. \((\color{red}{x}, \color{green}{y})\) \((\color{red}{0}, \color{green}{0})\) \((\color{red}{5}, \color{green}{6})\)

OpenStudy (mathstudent55):

Notice the points above. The x-coordinates are in red, and the y-coordinates are in green.

OpenStudy (mathstudent55):

\(slope = m = \dfrac{change ~in ~y}{change ~in ~x} \)

OpenStudy (mathstudent55):

The point (0, 0) has a y-coordinate of 0. The point (5, 6) has a y-coordinate of 6. How much did the y-value change from 0 to 6? What is 6 - 0 = ?

OpenStudy (anonymous):

6

OpenStudy (mathstudent55):

Correct. We now have this: \(slope = m = \dfrac{\color{green}{6}}{change ~in ~x}\)

OpenStudy (mathstudent55):

Now we do it for x.

OpenStudy (mathstudent55):

The point (0, 0) has an x-coordinate of 0. The point (5, 6) has an x-coordinate of 5. How much did the x-value change from 0 to 5? What is 5 - 0 = ?

OpenStudy (anonymous):

5

OpenStudy (mathstudent55):

Correct. We now have this: \(slope = m = \dfrac{\color{green}{6}}{\color{red}{5}}\)

OpenStudy (mathstudent55):

Since we now know the slope, and that was all that was missing in our equation, we can now finish this problem. The equation is \(y = mx\) \(y = \dfrac{6}{5}x\)

OpenStudy (anonymous):

Thank you SO much! @mathstudent55

OpenStudy (anonymous):

I got an 80% on my assessment! I passed it now I have 5 more left to do haha.

OpenStudy (mathstudent55):

You're welcome.